OpenStudy (anonymous):

Find the slope and the y intercept f(x)=-9x-10 it is an ordered pair... I know that x=o y=??

OpenStudy (anonymous):

to find y intercept you plug x = 0 and get f(0) = -10 so (0,-10) is the y intercept to find x intercept you replace f(x) by 0 (means y = 0) 0 = -9x - 10 9x = -10 x = -10/9 so (-10/9,0) is the x intercept

OpenStudy (anonymous):

so (0,9) ??

OpenStudy (anonymous):

what 0,9 ?

OpenStudy (anonymous):

never mind i missread sorry

OpenStudy (anonymous):

okay so what is the slope?

OpenStudy (anonymous):

y=f(x)=-9x-10 slope=-9 y-intercept=-10

OpenStudy (anonymous):

is the slope 9?

OpenStudy (anonymous):

-9x see the general line equation is : f(x) = mx + n m - slope n - y intercept

OpenStudy (anonymous):

ahh thanks :)

OpenStudy (anonymous):

no wonder i was thinking y=mx+b... blah

OpenStudy (anonymous):

slope=-9

OpenStudy (anonymous):

its the same

OpenStudy (anonymous):

oh it is?

OpenStudy (anonymous):

y = mx +b m - slope b - y intercept

OpenStudy (anonymous):

wow I see how well my brain is working today...lol

OpenStudy (anonymous):

instead of y they called it f(x) its a way to emphasize that it is a function of x

OpenStudy (anonymous):

@flycherrypie yes ur mind is brain is working fine